Studies in Bayesian Statistics

Workshop 7 - 2003

Friday, September 12 12:00 and 12:30 pm Bus leaves Holiday Inn for University Center 12:00 – 2:00 pm Registration - outside Connan Room 12:00 – 1:00 pm Poster presenters should assemble presentations in the

Dowd Room

1:00 – 4:00 pm Invited Case Study I McConomy Auditorium “Modeling Species Diversity Through Species Level Hierarchical Modeling” Authors: Alan E. Gelfand, John A. Silander, Jr., Shanshan Wu, Andrew Latimer and Paul Lewis Discussants: Jennifer Hoeting, Jay VerHoef

1:00 – 2:00 Presentation on background, description of problem, and role of statistical methods

2:00 – 2:15 Discussion

2:15 – 2:45 Presentation of statistical work

2:45 – 3:00 Break – Connan Room

3:00 – 3:30 Prepared discussion

3:30 – 4:00 Floor discussion and rejoinder

4:15 – 5:30 pm Seventh Morris H. DeGroot Memorial Lecture Guest Lecturer: Stephen M. Stigler, University of Chicago

5:30 – 6:30 pm DeGroot Lecture Reception - Ballroom 3 5:30 – 6:00 pm Poster presenters should use this time to assemble and/or

move their posters to the McKenna/Peter/Wright rooms.

6:00 – 7:30 pm Contributed Paper Poster Session McKenna/Peter /Wright rooms (cash bar)

7:30 – 9:00 pm Conference Dinner - Ballroom 1 and 2 9:00 – 10:30 pm Contributed Paper Poster Session McKenna/Peter/

Wright rooms (cash bar)

10:15 and 10:45 pm Bus departs University Center for Holiday Inn

Saturday, September 13 7:45 and 8:15 am Bus leaves Holiday Inn for University Center 8:00 – 8:30 am Continental Breakfast - Connan Room 8:30 – 11:30 am Invited Case Study II

McConomy Auditorium

“Highly Structured Models for Spectral and Image Analysis in High Energy Astrophysics” Authors: David A. van Dyk, David N. Esch, Yaming Yu, Margarita Karovska, Vinay L. Kashyap, Peter Freeman, Aneta Siemiginowska and Alanna Connors Discussants: Chris Genovese and Larry Wasserman, Ji-Meng Loh and Andrew Gelman

8:30 – 9:30 Presentation on background, description of problem, and role of statistical methods

9:30 – 9:45 Discussion

9:45 – 10:15 Presentation of statistical work

10:15 – 10:30 Break – Connan Room 10:30 – 11:00 Prepared discussion 11:00 – 11:30 Floor discussion and rejoinder

11:30 – 12:00 pm Lunch - Schatz Dining Room 11:45 – 1:00 pm Contributed Paper Poster Session - Ballroom 2-3 1:00 – 2:45 pm Panel discussion on Bioinformatics

McConomy Auditorium

Panel: Giovanni Parmigiani, John Storey, Mike West

2:45 – 3:00 pm Break – Connan Room 1:00 – 2:45 pm Panel discussion on Neuroinformatics McConomy Auditorium Panel: Emery Brown, Anthony Brockwell and Valerie Ventura

5:00 – 6:15 pm Refreshments and cash bar – Connan Room 5:15 – 6:15 pm Closing discussion - Connan Room 6:30 and 7:00 pm Bus departs University Center for Holiday Inn